OpenStudy (anonymous):

Which expression gives the solutions of –5 + 2x² = –6x?

myininaya (myininaya):

add 6x to both sides

myininaya (myininaya):

\[2x^2+6x-5=0\]

myininaya (myininaya):

Now use the quadratic formula! :)

myininaya (myininaya):

a=2,b=6,c=-5

OpenStudy (anonymous):

whats the quadratic formula?

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

i dont know how to use that. D:

myininaya (myininaya):

Do you know how to complete the square?

OpenStudy (anonymous):

no. :/ im horrible at math.

myininaya (myininaya):

hmmm... so what method do you want to use?

OpenStudy (anonymous):

i dont know any methods to solve that.

OpenStudy (ash2326):

@Qudrex I'll help you.

OpenStudy (anonymous):

okay.

OpenStudy (ash2326):

A standard quadratic equation is given as \[ax^2+bx+c=0\] its roots or its solution is given by the following formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (ash2326):

This formula gives us two values of x which satisfies the quadratic equation

OpenStudy (ash2326):

Did you understand till now?

OpenStudy (anonymous):

formulas confuse me.

OpenStudy (ash2326):

what's confusing you?

OpenStudy (anonymous):

the formula i dont know how to use it.

OpenStudy (ash2326):

It's very easy, I'll show you your question is -5+2x^2=6x The first step is to bring all the terms to one side \[-5+2x^2=6x\] Let's subtract both sides by 6x we get \[-5+2x^2-6x=6x-6x\] we'll get \[-5+2x^2-6x=0\] Now let's write this in descending powers of x \[2x^2-6x-5\] Did you understand till now?

OpenStudy (anonymous):

yes. i think i understand that.

OpenStudy (ash2326):

Great now let's compare this with standard equation \[ax^2+bx+c=0\] \[2x^2-6x+5=0\] we have \[a=2,\ b=-6\ and\ c=5 \] Did you get this?

OpenStudy (anonymous):

kinda

OpenStudy (ash2326):

Now its solution is given by \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] Let's substitute a=2, b=-6 and c=-5 \[x=\frac{-(-6)\pm \sqrt{(-6)^2-4\times (2)\times (-5)}}{2\times (2)}\] We get now \[x=\frac{6\pm \sqrt{36-(-40)}}{4}\] \[x=\frac{6\pm \sqrt{36+40}}{4}\] \[x=\frac{6\pm \sqrt{76}}{4}\] Did you understand this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

We've almost found the solution, we could simplify it more We have \[x=\frac{6\pm \sqrt{76}}{4}\] \[\sqrt {76}= \sqrt{4\times 19}=2\sqrt {19}\] so \[x=\frac{6\pm 2\sqrt{19}}{4}\] divide the whole equation by 2 we get \[x=\frac{3\pm \sqrt {19}}{2}\] so the two solutions\roots are \[x=\frac{3+ \sqrt {19}}{2}, \frac{3- \sqrt {19}}{2}\]\] Did you understand?

OpenStudy (anonymous):

yeah.

OpenStudy (ash2326):

So whenever you've to find roots of quadratic equation, use this formula:D

OpenStudy (anonymous):

okay!